OpenStudy (anonymous):

write a appropriate direct variation equation if y=27 when x=9

OpenStudy (solomonzelman):

The form is y=kx k is the constant of the direct variation. so, to find k, plug in numbers (and this is what the equation would be) 27=9k so we know that k=1/3

OpenStudy (anonymous):

thank you...

OpenStudy (solomonzelman):

You welcome, use this site it's good!

OpenStudy (anonymous):

Can you help me with this problem? Write an appropriate inverse variation equation if y-2 when x=9

OpenStudy (solomonzelman):

y=2? not y-2, correct?

OpenStudy (anonymous):

correct

OpenStudy (solomonzelman):

inverse variation form is y=k / x so if x=9 and y=2, then 2=k/9 the reason that y=k/x is called an inverse variation, is because the bigger the y the smaller the x, and vica versa. in your problem, note that the k=18 (as 9 times 2)

OpenStudy (anonymous):

Thank you...

OpenStudy (solomonzelman):

You welcome! \[\huge\color{blue}{ ✓ }\]